OpenStudy (anonymous):

Type in your answer as a reduced mixed number with the negative in front of the whole number. -8 3/4 + -11 1/3 =

OpenStudy (anonymous):

-20 1/12

OpenStudy (anonymous):

thanks!! :-)

OpenStudy (anonymous):

convert both numbers into full fractions 8*4 = 32 +3 = 35/4 11*3 = 33 + 1 = 34/3 common denominator of 4 so multiple the first by 3 3/3 * 35/4 = 105 / 12 second by 4 4/4 * 34/3 = 136 / 12 -105/12 + -136/12 = -241/12 = -20 and 1/12

OpenStudy (anonymous):

If you need to show work ^

OpenStudy (anonymous):

thanks soo much!! :-)